OpenStudy (anonymous):

Why is the point-slope form useful if you want to write the equation of a line through a point that is parallel to a given line.

OpenStudy (whpalmer4):

Because you'll need to make a line with the same slope for them to be parallel. Given the slope, and given a point, you're all set. point-slope form is y = mx + b where m is slope and b is y intercept. You'll have y, x, and m and you'll just need to plug in the known point to find b. For example, if the slope = 3, and the known point is (6,2), the line is y = 3x + b and you know that 2 = 3(6) + b so b must = -16 and the line is y = 3x - 16.